OpenStudy (anonymous):

Multiply or divide as indicated. x^2 * x^5

OpenStudy (hba):

Let's see this. When you multiply them you add the powers. x^a*x^b=x^a+b

OpenStudy (anonymous):

i dont get it sorry can you explain some more?

OpenStudy (hba):

Um let me give you the whole think. See the bases are the same. Let take another example a^3*a^3=a^3+3=a^6

OpenStudy (anonymous):

ok i think i get it thanks

OpenStudy (hba):

no problem :)

OpenStudy (anonymous):

so my answer would be x^7?

OpenStudy (anonymous):

Yes. That answer is correct.

OpenStudy (anonymous):

ok thanks

OpenStudy (hba):

Easy right? :D

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

can you help me with this one? Multiply or divide as indicated. (a+b)^9/(a+b)^4

OpenStudy (anonymous):

nvm i got it